What is the electric flux through each face of a cuboid?

Grade:12

3 Answers

bharat bajaj IIT Delhi
9 years ago
Now consider a cube... Charge inside the cube is 'q'. ..Total surface are of cube is 6*a*a(a is length of edge of cube). ..The total flux through the cube is q/epsilon...
q/epsilon is the flux for 6 faces...Therefore for one face the flux will be q/6epsilon..
Thanks
Bharat BajajIIT Delhi
askiitians faculty
Mostafijur Rahaman
9 years ago
i asked of a cuboid.
Aditya krishna
2 years ago
For applying Gauss’s law let we consider a cube formed by 2 cuboids.
-->now toatal flux(on cube) = q/E0.
-->total flux on 1 cuboid = q/2E0
→ Since cuboid has 6 faces 
→ 6 * (flux on one face)=q/(2E0).
Therefore flux on one side = q/12E0